# Efficient 6-dimensional phase space reconstruction from experimental measurements using generative machine learning

### v1

@@ IV Synthetic reconstruction Examples
In this section we evaluate the effectiveness of the GPSR technique on several different phase space distributions. For each case we describe below, the following procedure was followed. The first three quadrupoles (Q1, Q2, Q3) are tuned using Bayesian optimization implemented in Xopt [41], to minimize the transverse beam size at YAG2 with the scanning quadrupole (Q4) turned off. Once the range of quadrupole strengths is determined, we repeat a quadrupole scan four times, once for each combination of TDC and dipole states. Images from the quadrupole scans are then used to reconstruct the beam distribution. In the synthetic case studies that follow, the quadrupole strength was scanned over five steps, resulting in a total data set of 20 images.
@@ IV.1 Case 1: Gaussian beam reconstruction
The first case we explore is one where the beam distribution is a Multivariate Normal distribution that contains cross correlations between a number of the 6-dimensional phase space coordinates. Simulated measurements of the beam on the two diagnostic screens for each of the quadrupole scans is shown in Fig. 2. These images were then used to reconstruct the beam distribution using the GPSR approach with 100k macro particles. To maximize reconstruction accuracy, the reconstruction was trained for 3000 iterations of gradient descent which took roughly 10-15 minutes on an NVIDIA A100 GPU, although its possible that fewer iterations could be used.
The reconstruction algorithm results in a trained neural network that generates a distribution of macro-particles that should approximate the true beam distribution by matching simulated predictions to image measurements. The reconstructed beam distribution, with a comparison to the synthetic ground truth distribution, is shown in Fig. 3. We see from the 50th and 90th percentile contours (calculated by measuring the beam intensities that account for 50/90 percent of the total beam distribution) that the reconstructed beam distribution closely matches the true synthetic distribution. Furthermore, we can also compare measurements of the second order moments of the beam distribution, which also show close agreement between the reconstruction and the ground truth. In most cases the fractional error of the reconstructed second order moments is below the 10% level. In cases where the cross-covariances of the ground truth distribution are much smaller than the principal axis covariance the error of the reconstruction is higher, due to the dominance of the beam size along the principal axis over other beam features.
@@ IV.2 Case 2: Nonlinear beam reconstruction
We now analyze a case where the beam has a variety of non-linearities and correlations throughout the phase space distribution, as shown in Fig. 4 (simulated measurements of this beam distribution can be found in Appendix A). The reconstruction successfully replicates the ground truth distribution generated by performing 4 quadrupole scans with different states of the longitudinal phase space diagnostic elements, including non-linearities contained in cross-correlations of the beam distribution. Looking at the comparison between the predicted and ground truth values of the covariance matrix again shows good agreement within 10% for most elements.
Figure 4: Reconstruction results from non-linear synthetic beam distribution. Lower left: Comparison between 2-dimensional projections of the ground truth synthetic beam distribution with the reconstructed beam distribution. Solid lines denote ground truth projections and contours while dashed lines denote reconstruction predictions. White and grey contours denote 50th and 90th percentile intensity levels. Color map intensity denotes reconstructed prediction. Upper right: Comparison between ground truth and reconstructed second order moments of 90th percentile beam particles, where regular text denotes ground truth values and bold text denotes reconstructed predictions.
@@ IV.3 Case 3: EEX beam reconstruction
Detailed high-dimensional phase space reconstructions are particularly important when performing complex beam manipulations that impart significant correlations in 6-dimensional phase space distribution and contain precisely shaped features. An example of this is transverse-to-longitudinal emittance exchange (EEX) [42]. This process combines two doglegs and a transverse deflecting cavity to map the horizontal phase space distribution into a longitudinal distribution, allowing longitudinal profile shaping using transverse masking. For example, the EEX beamline has been used to generate ramped current profiles with a sharp drop off to improve the transformer ratio of dielectric [39] and plasma [43] wakefield acceleration. Precisely characterizing beam distributions created by the EEX beamline enables us to control accelerator and beam parameters, such as the location of leafs in a multi-leaf masking element [44] or focusing magnets before the EEX beamline [42], that lead to improved 6-dimensional beam tailoring for specific accelerator applications like wakefield accelerators.
@@
Reconstruction results from simulated diagnostic measurements (seen in Appendix A) of this beam distribution are shown in Fig. 5. We see that the reconstruction algorithm can resolve important features of the beam distribution, including the longitudinal profile and correlations between transverse and longitudinal phase spaces. However, the reconstruction algorithm has difficulty identifying large areas of uniform density within the beam, most notably in the $z-y$ phase space where the triangular head and the lower rectangular regions should both have near uniform density profiles. This leads to inaccurate predictions of the longitudinal beam profile, which is critical for high transformer ratio applications. Identifying how to improve the accuracy of the reconstruction in this case, either through algorithm modifications or changes in the diagnostic beamline, is a topic of future study.
@@ V Experimental Demonstration
In addition to simulation studies, we conducted an experimental study of using the 6-dimensional GPSR algorithm to characterize electron beams generated at AWA. We configured the AWA drive beamline to produce 1 nC electron bunches at a 2 Hz repetition rate with an energy of 43 MeV and transported them to the diagnostic configuration shown in Fig. 1 at the end of the AWA beamline. The beam charge was selected to provide a strong enough signal-to-noise ratio of the beam imaging diagnostics while mitigating coherent synchrotron radiation effects in the dipole spectrometer.
@@
Quadrupoles Q1-3 were then used to focus the beam onto YAG1 using Bayesian optimization algorithms while the scanning quadrupole Q4 was turned off. We then repeated 4 quadrupole scans (-2.9 T/m to 2.9 T/m, 9 steps) with different TDC and dipole settings as was done in the simulated examples. A set of 5 beam image shots were taken for each parameter setting with a charge window of 0.1 nC. The four parameter scans took approximately 8 minutes to perform at AWA, although we estimate that given better charge stability, the scan time could be reduced by a factor of 2.
@@
An additional complexity of the experimental measurement was a slight difference in image resolution between the two YAG screens used to measure the beam profile. These differences were incorporated into the reconstruction by defining two beamline simulations; one for the case where the dipole was off, and one where the dipole was on, each with different definitions for the screen diagnostic. The images were cropped to a size of 300 x 300 pixels, which corresponds to a side length of approximately 13 mm. Finally,the intensity of the image was clipped and set to zero at a lower bound threshold and a Gaussian smoothing filter was applied to the images to remove salt and pepper noise.
@@
To mitigate shot-to-shot jitter of the beam centroid, we shifted the distribution of each image such that the beam centroid was at the center of the region of interest and averaged the intensity profile over the 5 shots. As a result, our reconstruction predicts the structure of the distribution, but not its offset with respect to magnetic element centers, mean energy, or timing relative to the zero-crossing of the TDC cavity. A more robust treatment of the shot-to-shot jitter would allow us to identify these aspects of the beam distribution, as was done in [21], and is a topic of future study.
@@
In order to validate the accuracy of the reconstruction in the absence of a ground truth beam distribution, we compared model predictions to a subset of data that was not included in determining the phase space distribution (refereed to here as the test data set). If the generative model can accurately predict measurements inside the test set, we can have confidence that the reconstructed distribution is accurate. We selected every other quadrupole strength in the 4 quadrupole scans to be test data, resulting in a training data set consisting of 20 images (5 quadrupole settings x 4 LPS diagnostic settings) and a test data set consisting of 16 images (4 quadrupole settings x 4 LPS diagnostic settings). The entire data set can be viewed in Appendix A. With this training data set the reconstruction took approximately 17 minutes on an A100 Nvidia GPU.
@@
A comparison between predictions from the reconstruction and the experimentally measured test data is shown in Fig. 6. We see that the reconstruction accurately reproduces the 1D and 2D beam structure (with minor discrepancies in some cases) seen in the test images, including non-linearities and beamlet-like features in some cases. The reconstructed 6-dimensional phase space from experimental measurements is shown in Fig. 7. The reconstruction predicts normalized transverse emittances of $\varepsilon_{x,n}=19$ mm-mrad, $\varepsilon_{y,n}=8.5$ mm-mrad which are consistent with non-optimized AWA beamline parameters, the bunch charge of 1 nC, and the observed asymmetric beam size growth as a function of quadrupole strength when the dipole and deflecting cavity are off. One prominent feature of the distribution is the beamlet structure observed in the longitudinal current profile. We believe that this structure results from a set of 5 alpha-BBO crystals used at AWA to produce a longer, flat-top laser pulse profiles from 32 sub-pulses [45], however it is unclear how this leads to variations in beam energy along the bunch.
@@ VI Discussion
While the examples shown here demonstrate the effectiveness of the GPSR algorithm, the reliability and robustness of the algorithm can be improved for practical applications. The most critical next step for this work is to enable the algorithm to quantify uncertainties in the reconstructed phase space distribution from limited experimental measurements, especially in the context of phase space tomography. Conventional wisdom and empirical evidence [46] suggests that evenly-spaced angular rotations of the phase space over 360 degrees is necessary to accurately reconstruct 2-dimensional phase space distributions from 1-dimensional projections. However, it is not clear how this rule-of-thumb translates to reconstructing beam distributions in 6-dimensional phase space from 2-dimensional images, considering successes demonstrated here in accurately reconstructing different beam distributions using limited quadrupole scans, and considering nonlinear transformations inside the diagnostics used here, such as chromatic focusing effects. Developing an understanding of reconstruction uncertainty would help explain why a given collection of measurements would be sufficient to reconstruct a particular beam distribution. Additionally, this would aid in further reducing the number of measurements needed to reconstruct a particular beam distribution by tracking the reduction in reconstruction uncertainty as measurements are added.
@@
Finally, one advantage of GPSR is that the generative model creates these macro-particles by transforming samples from a simple random distribution, as a result, predicted particle distributions can contain any number of macro-particles regardless of the number of particles used during training the model. For example, even though the generative model in the cases demonstrated here was trained with 100k particles, reconstructions shown here predicts distributions containing one million macro-particles. This method for representing beam distributions is advantageous in a number of ways for use in simulations. Saving and transporting high-fidelity beam distributions containing a large number of 6-dimensional macro-particle coordinates is a memory intensive process and is restricted to a single-fidelity. Generative models on the other hand use only a few thousand scalar quantities (the weights and offsets of neural network parameters) and a description of the model structure to represent the beam distribution at any level of fidelity.
@@ VII Conclusion
In this work, we have demonstrated in several simulated and experimental case studies, that detailed characterizations of 6-dimensional beam distributions can be achieved rapidly using the GPSR algorithm such that it can be used in practical accelerator applications. This analysis method leverages the detail contained in 2-dimensional screen images of the beam distribution and knowledge of beam dynamics in the accelerator to significantly reduce (by up to a factor of 75x) the time needed to produce detailed predictions of the 6-dimensional phase space distribution. Furthermore, the generative ML representation of the beam distribution efficiently captures the complex structure of the beam distribution in the few thousand parameters of a small neural network, which then can be used to generate macro-particle distributions which can then be used for analysis or beam dynamics simulations. This technique has major implications for allowing 6-dimensional information to be used to inform accelerator control and understand complex physical phenomena.
@@ Acknowledgements.
Conceptualization, R.R., J.P.G.A., and A.L.E.; Data curation, R.R., and J.P.G.A.,; Formal analysis, R.R., and J.P.G.A.; Funding acquisition, A.L.E., J.P. and Y.K.K.; Investigation, R.R. and J.P.G.A.; Methodology, R.R., T.B., J.L.M., and A.L.E.; Software, R.R., and J.P.G.A.; Experiment, R.R., J.P.G.A., W.L., E.W., and A.O.; Supervision, A.L.E., R.R.,Y.K.K. and J.P.; Validation, R.R., and J.P.G.A.; Visualization, R.R., and J.P.G.A.; Writing—original draft, R.R., and J.P.G.A.; Writing—review and editing, R.R., J.P.G.A., and A.L.E. All authors have read and agreed to the published version of the manuscript. This work is supported by the U.S. Department of Energy, Office of Science under Contract No. DE-AC02-76SF00515 and the Center for Bright Beams, NSF award PHY-1549132. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231 using NERSC award ERCAP0020725.
@@ Appendix A Synthetic and Experimental Datasets
The following plots contain training data sets used to reconstruct the beam distributions described in Section IV and Section V .

### v2

@@ IV Synthetic reconstruction Examples
In this section, we evaluate the effectiveness of the GPSR technique on several different phase space distributions. For each case we describe below, the following procedure was followed. The first three quadrupoles (Q1, Q2, Q3) are tuned using Bayesian optimization implemented in Xopt [40] to minimize the transverse beam size at YAG1 with the scanning quadrupole (Q4) turned off. We then scan the focusing strength of Q4 four times, once for each combination of TDC and dipole states. Images from the quadrupole scans are then used to reconstruct the beam distribution. In the synthetic case studies that follow, the quadrupole strength was scanned over five steps, resulting in a total data set of 20 images.
@@ IV.1 Case 1: Gaussian beam reconstruction
The first case we explore is one where the beam distribution is a Multivariate Normal distribution that contains cross correlations between a number of the 6-dimensional phase space coordinates. Simulated measurements of the beam on the two diagnostic screens for each of the quadrupole scans is shown in Fig. 2. These images were then used to reconstruct the beam distribution using the GPSR approach with 100k macro particles. To maximize reconstruction accuracy, the reconstruction was trained for 3000 iterations of gradient descent (Adam [41], learning rate 0.01) which took roughly 10-15 minutes on an NVIDIA A100 GPU, although its possible that fewer iterations could be used.
The reconstruction algorithm results in a generative model which generates a distribution of macro-particles in 6-dimensional phase space that should approximate the true beam distribution. The reconstructed beam distribution, with a comparison to the synthetic ground truth distribution, is shown in Fig. 3. We see from the 50th and 90th percentile contours (calculated by measuring the beam intensities that account for 50/90 percent of the total beam distribution) that the reconstructed beam distribution closely matches the true synthetic distribution. Furthermore, we can also compare measurements of the second order moments of the beam distribution, which also show close agreement between the reconstruction and the ground truth. In most cases the fractional error of the reconstructed second order moments is below the 10% level. In cases where the cross-covariances of the ground truth distribution are much smaller than the principal axis covariances the error of the reconstruction is higher, due to the dominance of the beam size along the principal axis over other beam features.
@@ IV.2 Case 2: Nonlinear beam reconstruction
We now analyze a case where the beam has a variety of non-linearities and correlations throughout the phase space distribution, as shown in Fig. 4 (simulated measurements of this beam distribution can be found in Appendix A). The reconstruction successfully replicates the ground truth distribution generated by performing 4 quadrupole scans with different states of the longitudinal phase space diagnostic elements, including nonlinear features contained in cross-correlations of the beam distribution. Looking at the comparison between the predicted and ground truth values of the covariance matrix again shows good agreement within 10% for most elements.
Figure 4: Reconstruction results from nonlinear synthetic beam distribution. Lower left: Comparison between 2-dimensional projections of the ground truth synthetic beam distribution with the reconstructed beam distribution. Solid lines denote ground truth projections and contours while dashed lines denote reconstruction predictions. White and grey contours denote 50th and 90th percentile intensity levels. Color map intensity denotes reconstructed prediction. Upper right: Comparison between ground truth and reconstructed second order moments of 90th percentile beam particles, where regular text denotes ground truth values and bold text denotes reconstructed predictions.
@@ IV.3 Case 3: EEX beam reconstruction
Detailed 6-dimensional phase space reconstructions are particularly important when performing complex beam manipulations that impart significant correlations in the 6-dimensional phase space distribution and contain precisely shaped features. An example of this is transverse-to-longitudinal emittance exchange (EEX) [42, 25]. This process combines two doglegs and a transverse deflecting cavity to map the horizontal phase space distribution into a longitudinal distribution, allowing longitudinal profile shaping using transverse masking. For example, the EEX beamline has been used to generate ramped current profiles with a sharp drop off to improve the transformer ratio of dielectric [38] and plasma [43] wakefield acceleration. Precisely characterizing beam distributions created by the EEX beamline enables us to control accelerator and beam parameters, such as the location of leafs in a multi-leaf masking element [44] or focusing magnets before the EEX beamline [45], that lead to improved 6-dimensional beam tailoring for accelerator applications.
@@
Reconstruction results from simulated diagnostic measurements (seen in Appendix A) of this beam distribution are shown in Fig. 5. We see that the reconstruction algorithm can resolve important features of the beam distribution, including the longitudinal profile and correlations between transverse and longitudinal phase spaces. However, the reconstruction algorithm has difficulty identifying large areas of uniform density within the beam, most notably in the $z-y$ phase space where the triangular head and the lower rectangular regions should both have near uniform density profiles. This leads to slightly inaccurate predictions of the longitudinal beam profile, which is critical for high transformer ratio applications. Identifying how to improve the accuracy of the reconstruction in this case, either through algorithm modifications or changes in the diagnostic beamline, is a topic of future study.
@@ V Experimental Demonstration
In addition to simulation studies, we conducted an experimental demonstration of using the 6-dimensional GPSR algorithm to characterize electron beams generated at AWA. We configured the AWA drive beamline to produce 1 nC electron bunches at a 2 Hz repetition rate with an energy of 43 MeV and transported them to the diagnostic configuration shown in Fig. 1 at the end of the AWA beamline. The beam charge was selected to provide a strong enough signal-to-noise ratio for beam imaging diagnostics while mitigating coherent synchrotron radiation effects in the dipole spectrometer.
@@
Quadrupoles Q1-3 were then used to focus the beam onto YAG1 using Bayesian optimization algorithms while the scanning quadrupole Q4 was turned off. We then repeated 4 quadrupole scans (-2.9 T/m to 2.9 T/m, 9 steps) with different TDC and dipole settings, as was done in the simulated examples. A set of 5 beam image shots were taken for each parameter setting with a charge window of 0.1 nC. The four parameter scans took approximately 8 minutes to perform at AWA, although we estimate that given better charge stability, the scan time could be reduced by a factor of 2.
@@
An additional complexity of the experimental measurement was a slight difference in image resolution between the two YAG screens used to measure the beam profile due to minor differences in camera location and focusing. These differences were incorporated into the reconstruction by defining two beamline simulations; one for the case where the dipole was off, and one case where the dipole was on, each with different definitions for the screen diagnostic. The images were cropped to a square size of 300 x 300 pixels, which corresponds to a side length of approximately 13 mm. Finally, the intensity of the image was clipped and set to zero at a lower bound threshold and a Gaussian smoothing filter was applied to the images to remove salt and pepper noise.
@@
To mitigate shot-to-shot jitter of the beam, we shifted the distribution of each image such that the beam centroid was at the center of the region of interest and averaged the intensity profile over the 5 shots. As a result, our reconstruction predicts the structure of the distribution, but not its offset with respect to magnetic element centers, mean energy, or timing relative to the zero-crossing of the TDC cavity. A more robust treatment of the shot-to-shot jitter would allow us to identify these aspects of the beam distribution, as was done in [21], and is a topic of future study.
@@
In order to validate the accuracy of the reconstruction in the absence of a ground truth beam distribution, we compared model predictions to a subset of data that was not included in determining the phase space distribution (refereed to here as the test data set). If the generative model can accurately predict measurements inside the test set, we can have confidence that the reconstructed distribution is accurate. We selected every other quadrupole strength in the 4 quadrupole scans to be test data, resulting in a training data set consisting of 20 images (5 quadrupole settings x 4 LPS diagnostic settings) and a test data set consisting of 16 images (4 quadrupole settings x 4 LPS diagnostic settings). The entire data set can be viewed in Appendix A. With this training data set the reconstruction took approximately 17 minutes on an A100 Nvidia GPUs due to the increased image size compared to simulated examples.
@@
A comparison between predictions from the reconstruction and the experimentally measured test data is shown in Fig. 6. We see that the reconstruction accurately reproduces the 1 and 2-dimensional beam structure (with minor discrepancies in some cases) seen in the test images, including nonlinear and beamlet features in some cases. The reconstructed 6-dimensional phase space from experimental measurements is shown in Fig. 7. The reconstruction predicts normalized transverse emittances of $\varepsilon_{x,n}=19$ mm-mrad, $\varepsilon_{y,n}=8.5$ mm-mrad which are consistent with non-optimized AWA beamline parameters, the bunch charge of 1 nC, and the observed asymmetric beam size growth as a function of quadrupole strength when the dipole and deflecting cavity are off. One prominent feature of the distribution is the beamlet structure observed in the longitudinal current profile. We believe that this structure results from a set of 5 alpha-BBO crystals used at AWA to produce longer, flat-top laser pulse profiles by stacking 32 laser sub-pulses together [46], however it is unclear how this leads to variations in beam energy along the bunch.
@@ VI Discussion
While the examples shown here demonstrate the effectiveness of GPSR in these cases, the reliability and robustness of the algorithm can be improved for general applications. The most critical next step for this work is to enable the algorithm to quantify uncertainties in the reconstructed phase space distribution from limited experimental measurements. Conventional wisdom and empirical evidence [47] suggests that evenly-spaced angular rotations of the distribution over 360 degrees is necessary to accurately reconstruct 2-dimensional phase space distributions from 1-dimensional projections. However, it is not clear how this rule-of-thumb translates to reconstructing beam distributions in 6-dimensional phase space from 2-dimensional images or in the case where nonlinear beam manipulations are used. Developing an understanding of reconstruction uncertainty would help explain why a given collection of measurements would be sufficient to reconstruct a particular beam distribution. Additionally, this would aid in further reducing the number of measurements needed to reconstruct a particular beam distribution by tracking the reduction in reconstruction uncertainty as measurements are added to prevent redundant measurements.
@@
Finally, one advantage of GPSR is that the generative model creates these macro-particles by transforming samples from a simple random distribution. As a result, predicted particle distributions can contain any number of macro-particles regardless of the number of particles used during training the model. For example, even though the generative model in the cases demonstrated here was trained with 100k particles, the figures shown here uses predictive distributions containing one million macro-particles. This method for representing beam distributions is advantageous in a number of ways for use in simulations. Saving and transporting high-fidelity beam distributions containing a large number of 6-dimensional macro-particle coordinates is a memory intensive process. Generative models on the other hand use substantially fewer scalar quantities (the weights and offsets of neural network parameters) and a description of the model structure to represent the beam distribution at any level of fidelity, significantly reducing the memory needed to share the beam distribution between simulations.
@@ VII Conclusion
In this work, we have demonstrated in several simulated and experimental case studies that detailed characterizations of 6-dimensional beam distributions can be achieved rapidly using the GPSR algorithm. This analysis method leverages the detail contained in 2-dimensional screen images of the beam distribution and knowledge of beam dynamics in the accelerator to significantly reduce (by up to a factor of 75x) the time needed to produce detailed predictions of the 6-dimensional phase space distribution. Furthermore, the generative ML representation of the beam distribution is trained from scratch on experimental data, requiring no previous data collection or pre-training needed by other applications of ML in accelerator physics. As a result, the GPSR algorithm can be used to provide 6-dimensional phase space information during accelerator operations and in a wide variety of contexts. This technique has major implications for allowing 6-dimensional information to be used to inform accelerator control and understand complex physical phenomena.
@@ Acknowledgements.
Conceptualization, R.R., J.P.G.A., and A.L.E.; Data curation, R.R. and J.P.G.A.,; Formal analysis, R.R., and J.P.G.A.; Funding acquisition, A.L.E., J.P., and Y.K.K.; Investigation, R.R. and J.P.G.A.; Methodology, R.R., J.P.G.A., and A.L.E.; Software, R.R. and J.P.G.A.; Experiment, R.R., J.P.G.A., W.L., E.W., and A.O.; Supervision, A.L.E., R.R., Y.K.K., and J.P.; Validation, R.R., and J.P.G.A.; Visualization, R.R., and J.P.G.A.; Writing—original draft, R.R., and J.P.G.A.; Writing—review and editing, R.R., J.P.G.A., and A.L.E. All authors have read and agreed to the published version of the manuscript. This work is supported by the U.S. Department of Energy, Office of Science under Contract No. DE-AC02-76SF00515 and the Center for Bright Beams, NSF award PHY-1549132. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231 using NERSC award ERCAP0020725.
@@ Appendix A Synthetic and Experimental Datasets
Figures 89, and 10 contain training data sets used to reconstruct the beam distributions described in Section IV and Section V .